OpenStudy (tsuyu1):

Check my answer? A model rocket is launched from a roof into a large field. The path of the rocket can be modeled by the equation y = -0.02x^2 + 2.3x + 6, where x is the horizontal distance, in meters, from the starting point on the roof and y is the height, in meters, of the rocket above the ground. How far horizontally from its starting point will the rocket land? >57.50 m >115.00 m ***** >117.55 m >235.10 m ***** = my answer

OpenStudy (oops):

This is finding the zeros of the quadratic equation, and in this case, the simplest way to do that is the quadratic equation. \[\frac{ -2.3\pm \sqrt{2.3^2-4(-0.02)(6)} }{ 2(-0.02) }\] That all simplifies into \[\frac{ -2.3 \pm \sqrt{5.77} }{ -0.04 }\] But wait, there's more! That simplifies into some negative answer or 117.552060747, rounding to the 117.55. (I could be wrong but I would doublecheck yourself)